I dont know how to solve 5(x-3)+4=-3x+25=
Answer by sdmmadam@yahoo.com(530)

solve 5(x-3)+4=-3x+25
5(x-3)+4=-3x+25
(5x - 15)+4 =-3x+25
5x+(-15+4) =-3x +25 (by additive associativity)
5x-11 = -3x +25 (transposing,change side then change sign)
5x+3x = 25+11
8x = 36
x = 36/8 = (9X4)/(2X4) = 9/2
Answer x = 9/2
Verification: LHS =5(x-3)+4 = 5X(9/2-3)+4 = 5X(3/2)+4 = 15/2+4 = (15+8)/2 = 23/2
RHS = -3x+25 = (-3)X9/2 + 25 = -27/2 + 25 = 25-27/2 = (50-27)/2 = 23/2
LHS = RHS
Therefore our answer x = 9/2 is correct